### v1

# Radiant Triangle Soup with Soft Connectivity Forces for 3D Reconstruction and Novel View Synthesis

Abstract
In this work, we introduce an inference-time optimization framework utilizing triangles to represent the geometry and appearance of the scene. More specifically, we develop a scene optimization algorithm for triangle soup, a collection of disconnected semi-transparent triangle primitives. Compared to the current most-widely used primitives for 3D scene representation, namely Gaussian splats, triangles allow for more expressive color interpolation, and benefit from a large algorithmic infrastructure for downstream tasks. Triangles, unlike full-rank Gaussian kernels, naturally combine to form surfaces. We formulate connectivity forces between triangles during optimization, encouraging explicit, but soft, surface continuity in 3D. We perform experiments on a representative 3D reconstruction dataset and show competitive photometric and geometric results.
1 Introduction
Gaussian splatting has shown impressive results in Novel View Synthesis (NVS) and 3D reconstruction. However 3DGS [1] and the subsequent work fail to accurately model the geometry of surfaces. 3D Gaussians are by design diffuse volumetric primitives, which inherently causes difficulties in representing geometry and sharp boundaries. This is because they are optimized for novel view synthesis, which is achieved by alpha-blending several overlapping Gaussians, none of which has to be located on the underlying surfaces. In order to compensate, optimizing over these volumetric primitives leads to an expensive increase in the number of primitives and can cause floaters or blurry artifacts in the scene.
To rectify this issue, Gaussian Surfels [2, 3, 4] and General 3D Convexes [5] have been proposed to help with accuracy in geometry reconstruction. Gaussian Surfels use either strict 2D Gaussians [2] or flatten 3D Gaussians via loss regularization [3, 4] to better model thin surfaces, but these diffuse ellipsoids can still fail to accurately represent surfaces without substantially increasing their density. For clarity, throughout this paper we will refer to the work of Kerbl et al. [1] and Huang et al. [2] as 3DGS and 2DGS, respectively. We will use Volumetric Gaussian Splatting (VGS) and Surfel Gaussian Splatting (SGS) in general without referring to a particular approach.
In this paper, we present a new, complete methodology for 3D reconstruction that uses semi-transparent triangle primitives as the scene representation for Radiance Fields (RF). Our approach, which we named Radiant Triangle Soup (RTS), includes non-differentiable mechanisms for initialization, pruning, and densification, as well as differentiable mechanisms for rasterization. Our loss comprises 2D and 3D terms that favor more accurate rendering of color and surface normals as well as smoother surfaces. Within this framework, we have implemented an explicit Radiance Field algorithm and have achieved competitive results in terms of both appearance and geometry (See Fig. 1).
RTS provides a feature that is not supported by current GS scene representations, in which there is no avenue for explicit information sharing in 3D among primitives. Instead, Gaussian kernels communicate via alpha-blending when rendered onto common pixels. Back-propagating from the loss at each pixel to the primitives is the only means of coordination across Gaussians. Using multiple images from various viewpoints gives rise to several indirect constraints on the primitives, but direct constraints are currently absent. Our approach, on the other hand, is the first to enable direct information exchange among neighboring primitives. Having explicit 3D communication allows for a more direct coordination of primitive behavior throughout the optimization process.
Furthermore, most of the scene representations thus far operate with a single color parametrization for each primitive. This can lead to large numbers of overlapping primitives in order to reconstruct details in the surfaces of the scene. Using triangles as a representation allows for each triangle vertex to encode a separate color, leading to more expressive primitives through bilinear interpolation.
To summarize, the work presented here:
  • Develops a new SGS scene representation through alpha-blending of triangle primitives.
  • Introduces explicit 3D forces between primitives to encourage soft connectivity.
  • Incorporates a depth smoothness loss from the unsupervised Multi-View Stereo (MVS) literature.
Refer to caption Refer to caption Refer to caption
Figure 1: Optimizing over Radiant Triangle Soup (RTS) produces high quality 3D models from captured images. Above are the direct visualization of the point clouds without meshing from the DTU dataset [6]. See Section 6 for details.
2 Related Work
We begin this section with Gaussian splatting formulations that favor surfaces and continue with methods that rely on non-Gaussian primitives. Surveys on other aspects of Gaussian splatting, that we cannot cover due to space limitations, include [7, 8, 9, 10]
Gaussian Splatting
The seminal work of Kerbl et al. on 3D Gaussian Splatting [1] introduced an explicit alternative to NeRF [11] that is able to achieve high-quality rendering at much higher speed. 3DGS relies on interleaved differentiable optimization and non-differentiable density control of the explicit representation. The objective function of the former guides the optimization process to decrease view synthesis errors for one of the training images at each iteration.
2D Gaussian Splatting (2DGS) [2] modified 3DGS to prioritize the reconstruction of surfaces, rather than volumetric material. This was accomplished by collapsing the 3D GS to 2D discs by setting the minimum eigenvalue of the Gaussian to 0. Similar approaches were developed by the authors of SuGaR [3] and PGSR [4] who devised techniques for aligning 2D Gaussian with the surfaces via regularization and a multi-view loss, respectively. Gaussian Surfels [12] are also similarly relying on monocular surface normal estimates and a normal-depth consistency loss. Gaussian Opacity Fields (GOF) [13] enable direct geometry extraction as the zero-level set of 3D Gaussians without the need for external surface extraction modules.
3D-Half-Gaussian Splatting (3D-HGS) [14] enables the representation of perfectly planar surfaces and hard edges by attaching a splitting plane to each Gaussian and setting the density of one of the half-spaces to 0. The high-density splitting plane can then be aligned with the local surface.
Alternative Representations
Representations based on NeRF and 3DGS have found great success, but have also inspired researchers to seek alternatives. NeuRBF [15] utilizes Radial Basis Functions (RBFs) to overcome limitations of NeRF due to the global nature of its MLP and features. Instant-NGP [16] and Plenoxels [17] proposed representations with local features, but they were arranged in regular grids. In contrast, NeuRBF achieves spatial adaptivity and more precise handling of high frequency content via the use of anisotropic RBFs.
Recently, alternative primitives have begun to emerge. DARB-Splatting [18] is motivated by the observation that a broader class of decaying anisotropic RBFs (DARBFs), than just Gaussians, have 3D-to-2D projections that can be integrated easily. Compared to Gaussians, other DARBFs lead to faster convergence with a smaller memory footprint. GES [19] replaces the Gaussian kernel with a Generalized Exponential Function, overcoming the low-pass effect of the Gaussian and thus requiring far fewer primitives to represent the scene. Motivated by artifacts caused by the opacity of Gaussian splats falling off away from the centroid, SolidGS [20] modifies the Gaussian kernel by introducing a learnable solidness factor that makes the splats exhibit a more solid-like behavior. Chen et al. [21] argue that unlike Gaussian kernels that have unbounded support and may cause blurring, bounded linear kernels reduce foreground-background interference and yield clearer and more accurate reconstructions.
Besides the methods that force their Gaussians to be planar [2, 3, 4], but still diffuse, there are others that represent the scene with collections of planar primitives. Zanjani et al. [22] initially use Gaussian splats to model the scene and then merge them hierarchically to infer 3D planes, which are abundant in indoor scenes. PlanarSplatting [23] also addresses surface reconstruction in indoor scenes. In this case, planes are the only primitives, initialized via monocular depth estimation. TRIPS [24] uses points as primitives. The key principle is that points can be rasterized into an image pyramid from which the appropriate layer can be selected according to the size of the projected point. Holes can be filled by a small network yielding accurate, crisp renderings.
A few methods rely on volumetric primitives. LinPrim [25] is based on linear solid primitives, specifically octahedra and tetrahedra, with triangular faces that take advantage of well known fast processing and rendering mechanisms. LinPrim performs volumetric rendering by computing the entrance and exit points of each ray through the primitives. 3D Convex Splatting (3DCS) [5] was inspired by the same shortcoming of GS: very large numbers of primitives are needed to model hard edges and flat surfaces due to the diffuseness of the Gaussians. As the name suggests, the scene is represented by a set of polyhedral convexes which undergo volumetric rendering, pruning and splitting operations. The concepts of smoothness and sharpness introduced by 3DCS have inspired our diffuseness (see Section 4). Radiant Foam [26] is both efficient, like GS, but also enables modeling light transport phenomena like reflection and refraction. The method takes its name from the representation which is a dense Voronoi tessellation of 3D space. The positions of Voronoi sites are differentiable after rendering, allowing iterative optimization. Since this is the only representation with connected elements, non-differentiable operations, e.g. pruning, have to be implemented carefully. All methods in this paragraph rely on different forms of piecewise constant volumetric elements with planar faces. Computing the intersections of each primitive with a ray requires multiple ray-triangle intersections, and is thus more expensive than our method.
A few common themes emerge by analyzing the above methods. Most advocate the use of bounded primitives to enhance their ability to represent sharp edges and flat surfaces with small numbers of primitives. Our primitives are bounded but have diffuse boundaries to facilitate optimization. Like SolidGS [20], the opacity of our triangles remains approximately constant inside the triangle leading to sharper synthesized views. Several of the non-Gaussian primitives [5, 23, 25] leverage the existing software and hardware infrastructure by proposing representations with triangular faces.
3 Background
Before discussing our optimization framework, we first introduce the necessary background geometry used to parameterize a ray-triangle intersection in barycentric coordinates. The barycentric coordinates of a point on a triangle $\lambda=[\lambda^{0},\lambda^{1},\lambda^{2}]^{T}$ are one form of expressing the coordinates of the point as a linear combination of the coordinates of the three vertices, $[V^{0},V^{1},V^{2}]$. This parameterization is important for graphics applications to be able to efficiently rasterize triangles onto the screen space and interpolate the color from each vertex for a pixel. Throughout this paper, we follow the assumption that $\lambda^{j}\in\mathbb{R}^{+}$ corresponds to vertex $V^{j}\in\mathbb{R}^{3}$ such that $\sum_{j=1}^{3}\lambda^{j}=1$.
Given a pixel, we compute the ray-triangle intersection,
$P=C+(\widehat{r}d)$ (1)
where $C\in\mathbb{R}^{3}$ is the camera center, $\widehat{r}\in\mathbb{R}^{3}$ is the unit vector for the ray through pixel $i$, and $d\in\mathbb{R}$ is the depth along the ray from the camera center to the intersection point, computed as follows,
$d=\frac{\widehat{n}\cdot\overrightarrow{CB}}{\widehat{n}\cdot\widehat{r}}$ (2)
where $B\in\mathbb{R}^{3}$ is the barycenter of the triangle and $\widehat{n}\in\mathbb{R}^{3}$ is its normal.
We compute the barycentric coordinates for point $P$ using the triple products between the normal, a triangle edge, and the vector from each vertex to the point,
$\lambda=\frac{1}{\widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}V_{2}})}\begin{bmatrix}\widehat{n}\cdot(\overrightarrow{V_{1}V_{2}}\times\overrightarrow{V_{1}P})\\ \widehat{n}\cdot(\overrightarrow{V_{2}V_{0}}\times\overrightarrow{V_{2}P})\\ \widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}P})\end{bmatrix}$ (3)
Intuitively, the contribution of each vertex is proportional to the area of the sub-triangle formed by the intersection point $P$ and the other two vertices of the triangle. This weight becomes larger as $P$ approaches the vertex.
4 Method
Figure 2: We begin with a set of sparse points. From these points we initialize a set of triangles and compute nearest neighbor connections for each triangle edge. During optimization, we render an image, depth map, and normal map for each view. We compute both 2D loss (over all output renderings), as well as 3D loss (directly over the primitive connections). Adaptive densification is non-differentiable and is performed at set intervals throughout the optimization. Triangle edge neighbors are recomputed following adaptive densification.
Given a set of images together with camera poses and a set of sparse points $\mathcal{S}$ computed via Structure-From-Motion [27], we construct an explicit scene representation using triangle primitives, the parameterization of which we discuss in Section 4.1. To model the surfaces in the scene, we render into each camera the image, depth, and normal map of the triangles through alpha-blending (see Section 4.4). The triangle parameters are directly updated via back-propagation after computing losses between the rendering of the scene and the ground truth image from each view. We also include 2D losses on the rendered depth and normal maps (see Section 5). In our representation, primitives maintain soft connectivity with their neighbors, discussed in Section 4.5, for which we compute additional loss directly over the connection orientations (see Section 5). Similar to previous works [2, 3, 4, 1, 13], we develop an adaptive densification strategy in order to facilitate the addition and removal of primitives in the scene (see Section 4.6). We show an overview of our optimization algorithm in Fig. 2.
4.1 Parameterization
From the set $\mathcal{S}$ of sparse points, we first create an initial set of triangles $\mathcal{T}$. The triangle primitives $t_{n}\in\mathcal{T}$ in our scene representation are parametrized with,
$t_{n}=\{\mu,\Delta,s,R,\alpha,\sigma\}$ (4)
where $\mu$ is the triangle reference point, $\Delta=\begin{bmatrix}\delta^{0}&\delta^{1}&\delta^{2}\end{bmatrix}$ are the per-vertex Spherical Harmonics, $s=\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$ are the vertex scales, $R$ is the $3\times 3$ rotation matrix, $\alpha$ is the opacity, and $\sigma$ is the diffuse scalar, discussed in Section 4.3.
4.2 Initialization
The coordinates of the sparse SfM points are used as the initialization for the triangle reference points, and the point colors as the initialization for the zero-component of the Spherical Harmonics for all three vertices. Similar to previous works, the scale for our primitives is computed based on the average distance to the three nearest neighboring points. Each triangle starts as an equilateral triangle, using the scale to parameterize the distance of each vertex from the triangle reference point $\mu$ (see Fig. 3). Note that in this representation, the aforementioned triangle reference point $\mu$ is not the barycenter of the triangle; that is, $\mu$ is not guaranteed to be the mean of all three vertices, but instead represents the anchoring position of the triangle. The diffuse scalar $\sigma$ is also set as a function of the average distance to the three nearest neighboring points. The parameterization of the initial rotation matrix for each triangle is defined such that the $\overrightarrow{V^{2}V^{0}}$ edge and the triangle normal $\widehat{n}$ are used as the $x-$ and $z-$ axis, respectively. We then initialize each triangle by applying a random rotation.
Figure 3: Triangle parameterization. Each triangle is parameterized by a reference point $\mu$, and three scale values $\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$. These parameters, along with the rotation matrix $R$, define the coordinates of each vertex $V^{j}$.
4.3 Diffuse Primitive Boundaries
Alpha-blending is the primary driving force of optimization for all splatting frameworks. Blending across primitives facilitates optimization via gradient descent and encourages, among other behaviors, movement in primitive position and orientation. It is challenging in practice to optimize a scene with fully opaque primitives that do not smoothly blend with one another. It is therefore ideal to make the triangle primitives diffuse near the edges. Taking inspiration from 3DCS [5], we parameterize the response from each triangle as a function of the signed distance from the ray-triangle intersection to the nearest edge,
$w_{\sigma}=\frac{1}{1+e^{(\sigma l)}}$ (5)
where $w_{\sigma}$ is the diffuse weight, $l$ is the signed distance between the intersection and the nearest triangle edge in the plane of the triangle, and the scalar $\sigma$ is an optimizable parameter of each triangle that controls the level of diffuseness. As $\sigma$ increases, the boundary of the triangle becomes less diffuse, creating sharper primitive renderings (see Fig. 4). For the signed distance, $l<0$ occurs when the intersection point lies outside of the triangle boundary.
Refer to caption Refer to caption Refer to caption
$\longrightarrow$ increasing $\sigma$
Figure 4: Controlling triangle diffuseness with $\sigma$. Small values of $\sigma$ (left) result in diffuse triangles with extended range of influence. Large values of $\sigma$ (right) result in sharp triangles with no blurring across edges.
4.4 Rasterization
With the barycentric coordinates of the ray-triangles intersections and diffuse weight computed, we interpolate the color for the current pixel w.r.t. the colors of each vertex weighted by the barycentric coordinates,
$c_{n}=\begin{bmatrix}c^{0}&c^{1}&c^{2}\end{bmatrix}\lambda$ (6)
where $c^{0}$, $c^{1}$, and $c^{2}\in\mathbb{R}^{3}$ are computed from the SH components of each vertex, and $\lambda$ are the barycentric coordinates of the intersection point.
We aggregate the contribution to the current output pixel, $i$, from each intersected primitive,
$c_{i}=\sum w_{n}c_{n}$ (7)
where $w_{n}=\alpha w_{\sigma}T$, and $T=\Pi_{j=1}^{i-1}(1-\alpha_{j})$ is the transmittance for the current primitive [1].
Previous works [2, 4, 13] provide two methods for rendering per-pixel depth; (1) computing the average weighted intersection depth of all traversed primitives (mean depth), and (2) using only the depth of the primitive that causes the transmittance $T$ to exceed $0.5$ (median depth). In our work, we use median depth. Using the average weighted depth encourages the formation of many semi-transparent layers of primitives. Rendering median depth removes this blending and helps guide the formation of surfaces. We directly use $d$ from Eq. 2 for the depth of each intersection.
For each triangle, we compute the surface normal as the cross-product between two edges,
$\widehat{n}_{n}=\frac{\overrightarrow{V^{0}V^{2}}\times\overrightarrow{V^{0}V^{1}}}{||\overrightarrow{V^{0}V^{2}}\times\overrightarrow{V^{0}V^{1}}||_{2}}$ (8)
Unlike the rendered depth, to render per-pixel surface normals, we follow previous work [2, 4] and alpha-blend all the surface normals of all intersected primitive. The intuition is that through a normal consistency loss (see Section 5), the blended normals must align with the normals computed from the median depth map. This encourages all the intersected triangles to align (and eventually collapse onto) the median surface. Thus, we render per-pixel surface normals through alpha-blending,
$\overrightarrow{n}_{i}=\sum w_{n}\widehat{n}_{n}$ (9)
4.5 Connectivity
To enable soft connectivity among primitives and encourage flat, connected surfaces, we add connectivity and surface normal terms to the optimization objective. For each triangle edge, we assign, and periodically update, a connection to the nearest neighboring triangle edge. Taking inspiration from the energy functions introduced in 3D scene flow estimation [28], the loss of the connectivity increases according to the distance between their vertices and the inner product between their normals, discussed in Section 5. In order to prevent primitives from undesired connections, edge connection assignment must pass direction and orientation criteria, shown in Fig. 5. Simply, we must ensure that the neighboring triangle edge is in the correct direction and is oriented appropriately in order to connect with the current edge. We compute connectivity assignment and activate the connectivity losses after several epochs.
Refer to caption Refer to caption
(a) Direction Constraint (b) Orientation Constraint
Refer to caption Refer to caption
(c) $\tau$ (d) $\rho$
Figure 5: Above we show connectivity constraints between two triangles. For simplicity, it is assumed the triangles are co-planar.
Naively connecting with the nearest edge without any consideration of relative position may lead to connections that cross triangles and could cause them to flip. To prevent such connections, we compute three unit vectors: (1) the unit vector orthogonal to the current triangle edge and the current triangle normal (pointing outside the triangle), (2) the unit vector orthogonal to the neighboring triangle edge and the neighboring triangle normal (pointing outside the triangle), and (3) the connecting vector from the center of the current triangle edge to the center of the neighboring triangle edge (normalized to be unit length). To measure the relative position between the triangles, we take the inner product between these unit vectors. The direction criterion is measured as the inner product between the current triangle edge vector and the connecting vector, shown in Fig. 5(c). The orientation criterion is measured as the inner product between the current triangle edge vector and the neighboring triangle edge vector, shown in Fig. 5(d). We apply thresholds, $\rho$ and $\tau$ as shown in the figure, to these two values to prevent any unwanted connectivity from forming.
Intuitively, the direction criterion ensures the neighboring edge is in the correct direction w.r.t. the current triangle. Not satisfying this criterion would cause the current triangle to flip in order to reduce the connectivity loss. The orientation criterion ensures the edges are facing each other and the triangles have relatively similar surface normals. Not upholding this criterion would cause the neighboring triangle to flip in order to reduce the connectivity loss.
4.6 Adaptive Densification
Figure 6: During densification, large triangles are split into four sub-triangles with interpolated colors per-vertex.
We perform adaptive density control through the process of cloning and splitting triangles. Previous works [2, 1, 4, 3] perform the cloning and splitting procedures by duplicating primitives conditioned on scale and position gradients. In order to properly split large triangles, we must split them into four sub-triangles. During early iterations in the optimization, when triangles are split, the sub-triangles move independently to better align with surfaces in the scene. During the later stages when most of the surfaces have formed, the triangles split to enable more detailed surface color representation, remaining relatively attached via the connectivity forces. See Fig. 6 for a visualization of the splitting procedure. Similar to previous work, we directly clone small primitives selected for densification.
Much like 2DGS [2], our method does not directly rely on the gradient of the projected 2D primitive center. Instead of computing an approximation via projecting 3D gradients into the camera plane [2], we directly condition densification on the magnitude of the reference point gradient $\nabla\mu$ in 3D.
5 Optimization
Our objective function comprises terms computed in 2D, on the images planes, and in 3D, on the triangles.
5.1 Rendering Losses
Following previous work [1, 2, 3, 4, 13], we compute the SSIM between the rendered and input images,
$\mathcal{L}_{ssim}=\frac{1}{N}\sum_{i,j}(1-\gamma)L_{1}+\gamma L_{ssim}$ (10)
To help locally align the triangles with the rendered surface, we apply a normal consistency loss [2]. This encourages alignment between the rendered normal $\widehat{n}$ with the normal computed from the rendered depth map $\widehat{n}_{d}$,
$\mathcal{L}_{norm}=\frac{1}{N}\sum_{i,j}(1-\widehat{n}^{T}\widehat{n}_{d})$ (11)
From the unsupervised depth estimation literature [29, 30, 31, 32, 33], we adopt a smoothness term on the rendered depth map conditioned on the gradient of the input image. This penalizes large gradients in the rendered depth map where we have small gradients in the input image,
$\mathcal{L}_{smooth}=\frac{1}{N}\sum_{i,j}||\partial_{x}D_{i,j}||e^{||\partial_{x}I_{i,j}||}+||\partial_{y}D_{i,j}||e^{||\partial_{y}I_{i,j}||}$ (12)
5.2 Scene Loss
To encourage connectivity between primitives, we penalize the mean of the $L2$ distance between the connected vertex pairs of neighboring triangle edges,
$\mathcal{L}_{conn}=\sum_{a\in\Omega}\frac{1}{2}(||V_{a}^{1}-V_{b}^{1}||_{2}+||V_{a}^{2}-V_{b}^{2}||_{2})+(1-\widehat{n}_{a}^{T}\widehat{n}_{b})$ (13)
where $V_{a}^{j}$ and $V_{b}^{j}$ are the $j^{th}$ vertex pair of connected edges $a$ and $b$, respectively. $\widehat{n}_{a}$ and $\widehat{n}_{b}$ are the normals of the connecting triangles and $\Omega$ is the set of all triangles that intersect the current camera frustum. Activating this loss for invisible triangles without rendering losses leads to oversmoothing.
The final objective is a weighted summation of the above,
$\mathcal{L}=\omega_{0}\mathcal{L}_{ssim}+\omega_{1}\mathcal{L}_{norm}+\omega_{2}\mathcal{L}_{smooth}+\omega_{3}\mathcal{L}_{conn}$ (14)
6 Experiments
Table 1: Chamfer distance evaluation on scenes from DTU [6]. Following previous literature, we average the accuracy and completeness Chamfer distances on the widely used evaluation set. Chamfer distances are measured in millimeters.
Method DTU
24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean (mm)$\downarrow$
3DGS [1] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96
SuGaR [3] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
2DGS [2] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
GOF [13] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
PGSR [4] 0.36 0.57 0.38 0.33 0.78 0.58 0.50 1.08 0.63 0.59 0.46 0.54 0.30 0.38 0.34 0.52
RTS 0.48 0.59 0.79 0.39 0.58 0.92 0.77 0.94 0.77 0.37 0.79 1.12 0.37 0.71 0.49 0.67
6.1 Implementation Details
We implement the majority of our RTS framework in Python using PyTorch [34]. For rasterization, we develop custom CUDA kernels for both the forward and backward pass. We run all our experiments on a single NVIDIA RTX A6000. In all of our experiments, we choose $\tau=0.0$ and $\rho=0.0$ for the threshold values on the connectivity constraints. For our loss weights, we choose $\omega=\begin{bmatrix}1.0&0.05&0.8&10.0\end{bmatrix}$ empirically.
Following state-of-the-art Multi-View Stereo methods [35, 36, 37, 38, 39, 40], we directly generate a final 3D point cloud from the rendered depth maps. To generate our final point cloud, we use heuristic filtering, similar to the post-processing presented in GBiNet [40]. For each rendered depth map, we project each depth estimate into neighboring views, using the depth estimates in each view at the projected pixel location to reproject back to the reference view, measuring the pixel reprojection error and filtering out estimates whose error is above a threshold.
6.2 Evaluation
We test our framework on DTU [6], which is an indoor dataset that contains images of 124 scenes taken from a camera mounted on an industrial robot arm. All scenes share the same camera trajectories, with ground-truth point clouds captured via structured light.
We evaluate our new approach on the DTU dataset and record the Chamfer distance in Table 1. We show competitive results alongside the leading state-of-the-art SGS methods. Across all scenes in the test set, we report highly competitive metrics against the current state of the art. Notably, our method seems particularly sensitive to specific scenes with texture-less and non-Lambertian surfaces. Qualitative results can be seen in Fig. 7.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ground Truth Rendered Image Rendered Depth Rendered Normal Point Cloud
Figure 7: Qualitative results on several scenes from the DTU dataset [6].
We measure the quality of novel view synthesis for our method on the DTU dataset using standard metrics of PSNR, SSIM, and LPIPS [41] in Table 2. Our method produces comparable results, with a slight reduction in PSNR. As von Lützow and Nießner [25] note, using primitives with explicit boundaries can begin to introduce hard edges in regions with poor visibility. While the diffuse boundaries of our triangles helps prevent much of this behavior, its effects are noticeable in some of the reconstructed images.
Table 2: Photometric evaluation on several scenes from the DTU [6] dataset. We record the per-scene SSIM, PSNR, and LPIPS metrics on the widely used test set.
Method DTU
SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
SuGaR [3] - 34.57 -
2DGS [2] 0.939 34.33 0.169
PGSR [4] 0.917 33.18 0.251
RTS 0.918 31.68 0.213
7 Limitations & Conclusions
Our geometric loss terms, much like 2DGS, involves a trade-off between model reconstruction quality and novel-view synthesis. Adhering more strictly to the geometric losses in our framework can lead to over-smoothing in some areas, and at times can prohibit the movement of primitives, as triangles are more strongly attracted to each other, leaving parts of the scene under-reconstructed. Furthermore, due to the periodic nearest-neighbors search, the run-time for our method increases with the number of primitives. Because of this, it can become difficult to reconstruct very large scenes.
In this work, we introduce a new scene representation, namely Radiant Triangle Soup (RTS). To the best of our knowledge, we are the first to introduce explicit 3D forces between primitives in a splatting framework, helping to coordinate the positioning of primitives to directly form surfaces. The introduction of Triangle Soup as the scene representation for radiance fields is amenable to future work in surface optimization and can be augmented to include more extensive forms of triangle connectivity and mesh evolution. After some modifications, our framework can also be used for optimization over watertight meshes. In future work, we plan to extend RTS with modified primitive connectivity strategies and mesh post-processing refinement.
Supplementary Material
S.1 Implementation Details
In this section, we describe our experimental setting and optimization parameters in detail. To begin optimization, similar to previous work [4, 2], all geometric supervision is disabled, with optimization only being guided initially by the SSIM loss $\mathcal{L}_{ssim}$. We enable the normal consistency loss $\mathcal{L}_{norm}$ at iteration $7,000$ and enable both the smoothness loss $\mathcal{L}_{smooth}$ and connectivity loss $\mathcal{L}_{conn}$ at iteration $10,000$, both chosen empirically. All triangles start with an initial opacity ($\alpha$) value set to $0.1$, with opacity for all primitives being reset every $3,000$ iterations. We run optimization on the scenes from the DTU and Mip-NeRF 360 datasets for $25,000$ and $30,000$ iterations, respectively. Densification and pruning is run every $250$ iterations starting after iteration $2,000$. The maximum reference point gradient threshold for densification is set to $7.5e^{-5}$ in all experiments.
The learning rates for each respective parameter are set as follows:
  • Spherical Harmonics ($\Delta$): $2.5e^{-3}$
  • Opacity ($\alpha$): $5e^{-2}$
  • Reference Point ($\mu$): $\begin{bmatrix}1.5e^{-4},&2e^{-6}\end{bmatrix}$
  • Rotation ($R$): $1e^{-3}$
  • Scale ($s$): $4e^{-3}$
  • Diffuse Scalar ($\sigma$): $1e^{-3}$
where the Reference Point learning rate follows an exponential decay scheduler starting with $1.5e^{-4}$ and ending with $2e^{-6}$ conditioned on the total number of iterations. We also tested a linear decay scheduler leading to similar results.
The run-times for our approach are roughly $1.5$ hours on DTU and $4.5$ hours on Mip-NeRF 360, with scenes from DTU and scenes from Mip-NeRF 360 having on average $249,941$ and $1,275,985$ primitives, respectively, using $0.5$ resolution for the DTU dataset and $0.25$ resolution for the Mip-NeRF 360 dataset.
S.2 Ablation on Depth Rendering
We show an ablation on the effects of using mean and median depth, in Table S.1. While using median depth has virtually no effect on the quality of the reconstructed images, there is a substantial effect on the accuracy and consistency of the rendered depth maps, and thus, the final geometry of the 3D model. The normal consistency loss $\mathcal{L}_{norm}$ is indifferent to surface layering when computed w.r.t. the blended mean depth and the blended mean normal. Using the median depth guides the median triangle to better align with the blended normals and vice versa.
Table S.1: Comparison of the effects of mean vs median depth using the entire DTU evaluation set.
Method Acc.(mm) $\downarrow$ Comp.(mm) $\downarrow$ PSNR $\uparrow$ Primitives $\downarrow$
Mean Depth 1.23 1.36 31.69 228,000
Median Depth 0.64 0.70 31.68 249,000
Table S.2: Novel View Synthesis on all scenes from the Mip-NeRF 360 [42] dataset.
Method Outdoor Indoor
[…]
NeRF [11] 21.46 0.458 0.515 26.84 0.790 0.370
Deep Blending [43] 21.54 0.524 0.364 26.40 0.844 0.261
Instant NGP [16] 22.90 0.566 0.371 29.15 0.880 0.216
MipNeRF360 [42] 24.47 0.691 0.283 31.72 0.917 0.180
SuGaR [3] 22.93 0.629 0.356 29.43 0.906 0.225
3DGS [1] 24.64 0.731 0.234 30.41 0.920 0.189
2DGS [2] 24.34 0.717 0.246 30.40 0.916 0.195
GOF [13] 24.82 0.750 0.202 30.79 0.924 0.184
PGSR [4] 24.76 0.752 0.203 30.36 0.934 0.147
3DCS [5] 24.07 0.700 0.238 31.33 0.927 0.166
RTS 21.41 0.657 0.349 30.28 0.921 0.130
S.3 Additional Evaluations
In this subsection, we show additional results on the Mip-NeRF 360 dataset [42]. Following the protocol specified by the authors of Mip-NeRF 360, we separate the images in each scene, taking every eighth image as a test image and training on the remaining. As standard evaluation, we report PSNR, SSIM, and LPIPS [41] metrics. We show results on all indoor and outdoor scenes in Table S.2. For indoor scenes, RTS shows results on par with the state-of-the-art methods, having the leading LPIPS score, the fourth overall SSIM score, and a highly competitive PSNR score among all listed methods. On two outdoor scenes (treehill and flowers), RTS shows difficulty reconstructing extremely distant background foliage, impacting the overall metrics.
We provide supplemental renderings of novel views on challenging indoor and outdoor scenes with fine details and non-Lambertian surfaces. Output renderings and 3D models for all scenes will be made publicly available to the research community if the paper is accepted.
We show qualitative geometric reconstruction results on all scenes from the DTU dataset in Fig. S.2 and novel view synthesis results on all indoor and outdoor scenes from the Mip-NeRF 360 datasets in Fig. S.1 and Fig. S.3, respectively.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(a) GT Image (b) Rendered Image (c) Rendered Depth
Figure S.1: Visualizations of all indoor scenes, (top-down) [bonsai, counter, kitchen, room], from the Mip-NeRF 360 dataset [42].
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure S.2: Point cloud visualization for all scenes from the evaluation set from the DTU dataset [6].
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(a) GT Image (b) Rendered Image (c) Rendered Depth
Figure S.3: Visualizations of all outdoor scenes, (top-down) [bicycle, flowers, garden, stump, treehill], from the Mip-NeRF 360 dataset [42].

### v2

@@
 Abstract
-In this work, we introduce an inference-time optimization framework utilizing triangles to represent the geometry and appearance of the scene.
-More specifically, we develop a scene optimization algorithm for triangle soup, a collection of disconnected semi-transparent triangle primitives.
-Compared to the current most-widely used primitives for 3D scene representation, namely Gaussian splats, triangles allow for more expressive color interpolation, and benefit from a large algorithmic infrastructure for downstream tasks.
-Triangles, unlike full-rank Gaussian kernels, naturally combine to form surfaces.
-We formulate connectivity forces between triangles during optimization, encouraging explicit, but soft, surface continuity in 3D.
-We perform experiments on a representative 3D reconstruction dataset and show competitive photometric and geometric results.
+We introduce an inference-time scene optimization algorithm utilizing triangle soup, a collection of disconnected translucent triangle primitives, as the representation for the geometry and appearance of a scene.
+Unlike full-rank Gaussian kernels, triangles are a natural, locally-flat proxy for surfaces that can be connected to achieve highly complex geometry.
+When coupled with per-vertex Spherical Harmonics (SH), triangles provide a rich visual representation without incurring an expensive increase in primitives.
+We leverage our new representation to incorporate optimization objectives and enforce spatial regularization directly on the underlying primitives.
+The main differentiator of our approach is the definition and enforcement of soft connectivity forces between triangles during optimization, encouraging explicit, but soft, surface continuity in 3D.
+Experiments on representative 3D reconstruction and novel view synthesis datasets show improvements in geometric accuracy compared to current state-of-the-art algorithms without sacrificing visual fidelity.
+Figure 1: Optimizing Radiant Triangle Soup (RTS) produces high quality 3D models from captured images.
+Above, from left to right, are the rendered normal map, rendered depth map, rendered image, and the direct rasterization of the triangle primitives, as if they were opaque.
 1 Introduction
-Gaussian splatting has shown impressive results in Novel View Synthesis (NVS) and 3D reconstruction.
-However 3DGS [1] and the subsequent work fail to accurately model the geometry of surfaces.
-3D Gaussians are by design diffuse volumetric primitives, which inherently causes difficulties in representing geometry and sharp boundaries.
-This is because they are optimized for novel view synthesis, which is achieved by alpha-blending several overlapping Gaussians, none of which has to be located on the underlying surfaces.
-In order to compensate, optimizing over these volumetric primitives leads to an expensive increase in the number of primitives and can cause floaters or blurry artifacts in the scene.
-To rectify this issue, Gaussian Surfels [2, 3, 4] and General 3D Convexes [5] have been proposed to help with accuracy in geometry reconstruction.
-Gaussian Surfels use either strict 2D Gaussians [2] or flatten 3D Gaussians via loss regularization [3, 4] to better model thin surfaces, but these diffuse ellipsoids can still fail to accurately represent surfaces without substantially increasing their density.
-For clarity, throughout this paper we will refer to the work of Kerbl et al.
-[1] and Huang et al.
-[2] as 3DGS and 2DGS, respectively.
-We will use Volumetric Gaussian Splatting (VGS) and Surfel Gaussian Splatting (SGS) in general without referring to a particular approach.
-In this paper, we present a new, complete methodology for 3D reconstruction that uses semi-transparent triangle primitives as the scene representation for Radiance Fields (RF).
-Our approach, which we named Radiant Triangle Soup (RTS), includes non-differentiable mechanisms for initialization, pruning, and densification, as well as differentiable mechanisms for rasterization.
-Our loss comprises 2D and 3D terms that favor more accurate rendering of color and surface normals as well as smoother surfaces.
-Within this framework, we have implemented an explicit Radiance Field algorithm and have achieved competitive results in terms of both appearance and geometry (See Fig.
+Gaussian Splatting (GS) methods are effective at Novel View Synthesis (NVS).
+However 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) fails to accurately model the geometry of scene surfaces.
+3D Gaussians are by design smooth, unbounded volumetric primitives, which are inherently ill-suited for representing flat surfaces and sharp boundaries.
+Optimization for novel view synthesis involves alpha-blending several overlapping Gaussians, none of which have to be located on the underlying surfaces.
+As a result, optimizing a representation of 3D Gaussian kernels for image synthesis commonly causes floaters or blurry artifacts in the scene.
+Several authors (Huang et al., 2024; Guédon & Lepetit, 2024; Chen et al., 2024a; Dai et al., 2024) have proposed flattening the kernels, with either strict 2D Gaussians (Huang et al., 2024) or flattened 3D Gaussians via loss regularization (Guédon & Lepetit, 2024; Chen et al., 2024a), to better model thin surfaces.
+These methods have shown that modifying the underlying primitives and enforcing optimization objectives on rendered geometry leads to improvements in geometric accuracy.
+While flat primitive help with surface alignment, diffuse ellipsoids still fail to accurately model depth discontinuities (see Fig.
+8).
+Triangles are the fundamental primitives in computer graphics because they are necessarily planar and convex.
+They can approximate any surface as piecewise planar at any tolerance level, can directly model sharp discontinuities, and can be rendered rapidly without approximations.
+We present a new methodology for 3D reconstruction, which we named Radiant Triangle Soup (RTS) that enables gradient-based optimization of a Radiance Field (RF) using translucent triangle primitives as the scene representation (see Fig.
 1).
-RTS provides a feature that is not supported by current GS scene representations, in which there is no avenue for explicit information sharing in 3D among primitives.
-Instead, Gaussian kernels communicate via alpha-blending when rendered onto common pixels.
+We provide a complete framework, including differentiable mechanisms for rasterization, as well as non-differentiable mechanisms for initialization, pruning, and densification.
+Our experiments show RTS achieves competitive results in terms of both appearance and geometry (see Fig.
+7).
+Conceptually, RTS provides a feature that is not supported by any other GS scene representation: an avenue for explicit information sharing among 3D among primitives.
+Conventional Gaussian kernels interact via alpha-blending when rendered onto common pixels.
 Back-propagating from the loss at each pixel to the primitives is the only means of coordination across Gaussians.
 Using multiple images from various viewpoints gives rise to several indirect constraints on the primitives, but direct constraints are currently absent.
-Our approach, on the other hand, is the first to enable direct information exchange among neighboring primitives.
-Having explicit 3D communication allows for a more direct coordination of primitive behavior throughout the optimization process.
-Furthermore, most of the scene representations thus far operate with a single color parametrization for each primitive.
-This can lead to large numbers of overlapping primitives in order to reconstruct details in the surfaces of the scene.
-Using triangles as a representation allows for each triangle vertex to encode a separate color, leading to more expressive primitives through bilinear interpolation.
+Conversely, RTS is the first framework to enable direct information exchange among neighboring primitives.
+We formulate connectivity losses between neighboring triangle edges, allowing for a more direct and effective coordination of, and constraint on, primitive behavior throughout the optimization process.
+Even though using triangles as primitives alone achieves better geometric accuracy compared to Gaussian surfels, encouraging connectivity during optimization leads to more accurate surfaces with less floating artifacts (see Table 3).
+Color expressivity on a per-primitive basis is another dimension where the RTS representation excels.
+It is common practice to use a single color parametrization for each primitive.
+However, this may require large numbers of overlapping primitives in order to reconstruct details in the surfaces of the scene.
+Using triangles as the representation allows for each vertex to encode a separate color (in the form of Spherical Harmonics), leading to more expressive primitives through bilinear interpolation.
 To summarize, the work presented here:
-  • Develops a new SGS scene representation through alpha-blending of triangle primitives.
+  • Develops a new scene representation through alpha-blending of triangle primitives.
   • Introduces explicit 3D forces between primitives to encourage soft connectivity.
-  • Incorporates a depth smoothness loss from the unsupervised Multi-View Stereo (MVS) literature.
-Refer to caption Refer to caption Refer to caption
-Figure 1: Optimizing over Radiant Triangle Soup (RTS) produces high quality 3D models from captured images.
-Above are the direct visualization of the point clouds without meshing from the DTU dataset [6].
-See Section 6 for details.
+  • Increases primitive expressivity via multi-color encoding.
 2 Related Work
 We begin this section with Gaussian splatting formulations that favor surfaces and continue with methods that rely on non-Gaussian primitives.
-Surveys on other aspects of Gaussian splatting, that we cannot cover due to space limitations, include [7, 8, 9, 10]
-Gaussian Splatting
+Surveys on other aspects of Gaussian splatting, omitted due to space limitations, include (Bao et al., 2025; Dalal et al., 2024; Luo et al., 2024; Wu et al., 2024).
 The seminal work of Kerbl et al.
-on 3D Gaussian Splatting [1] introduced an explicit alternative to NeRF [11] that is able to achieve high-quality rendering at much higher speed.
+(2023) on 3D Gaussian Splatting introduced an explicit alternative to NeRF (Mildenhall et al., 2020) that is able to achieve high-quality rendering at much higher speed.
 3DGS relies on interleaved differentiable optimization and non-differentiable density control of the explicit representation.
-The objective function of the former guides the optimization process to decrease view synthesis errors for one of the training images at each iteration.
-2D Gaussian Splatting (2DGS) [2] modified 3DGS to prioritize the reconstruction of surfaces, rather than volumetric material.
+The optimization process decreases view synthesis errors for one of the training images at each iteration and density control guides the placement of primitives.
+2D Gaussian Splatting (2DGS) (Huang et al., 2024) modified 3DGS to prioritize the reconstruction of surfaces, rather than volumetric material.
 This was accomplished by collapsing the 3D GS to 2D discs by setting the minimum eigenvalue of the Gaussian to 0.
-Similar approaches were developed by the authors of SuGaR [3] and PGSR [4] who devised techniques for aligning 2D Gaussian with the surfaces via regularization and a multi-view loss, respectively.
-Gaussian Surfels [12] are also similarly relying on monocular surface normal estimates and a normal-depth consistency loss.
-Gaussian Opacity Fields (GOF) [13] enable direct geometry extraction as the zero-level set of 3D Gaussians without the need for external surface extraction modules.
-3D-Half-Gaussian Splatting (3D-HGS) [14] enables the representation of perfectly planar surfaces and hard edges by attaching a splitting plane to each Gaussian and setting the density of one of the half-spaces to 0.
-The high-density splitting plane can then be aligned with the local surface.
-Alternative Representations
-Representations based on NeRF and 3DGS have found great success, but have also inspired researchers to seek alternatives.
-NeuRBF [15] utilizes Radial Basis Functions (RBFs) to overcome limitations of NeRF due to the global nature of its MLP and features.
-Instant-NGP [16] and Plenoxels [17] proposed representations with local features, but they were arranged in regular grids.
-In contrast, NeuRBF achieves spatial adaptivity and more precise handling of high frequency content via the use of anisotropic RBFs.
-Recently, alternative primitives have begun to emerge.
-DARB-Splatting [18] is motivated by the observation that a broader class of decaying anisotropic RBFs (DARBFs), than just Gaussians, have 3D-to-2D projections that can be integrated easily.
-Compared to Gaussians, other DARBFs lead to faster convergence with a smaller memory footprint.
-GES [19] replaces the Gaussian kernel with a Generalized Exponential Function, overcoming the low-pass effect of the Gaussian and thus requiring far fewer primitives to represent the scene.
-Motivated by artifacts caused by the opacity of Gaussian splats falling off away from the centroid, SolidGS [20] modifies the Gaussian kernel by introducing a learnable solidness factor that makes the splats exhibit a more solid-like behavior.
-Chen et al.
-[21] argue that unlike Gaussian kernels that have unbounded support and may cause blurring, bounded linear kernels reduce foreground-background interference and yield clearer and more accurate reconstructions.
-Besides the methods that force their Gaussians to be planar [2, 3, 4], but still diffuse, there are others that represent the scene with collections of planar primitives.
-Zanjani et al.
-[22] initially use Gaussian splats to model the scene and then merge them hierarchically to infer 3D planes, which are abundant in indoor scenes.
-PlanarSplatting [23] also addresses surface reconstruction in indoor scenes.
-In this case, planes are the only primitives, initialized via monocular depth estimation.
-TRIPS [24] uses points as primitives.
-The key principle is that points can be rasterized into an image pyramid from which the appropriate layer can be selected according to the size of the projected point.
-Holes can be filled by a small network yielding accurate, crisp renderings.
-A few methods rely on volumetric primitives.
-LinPrim [25] is based on linear solid primitives, specifically octahedra and tetrahedra, with triangular faces that take advantage of well known fast processing and rendering mechanisms.
-LinPrim performs volumetric rendering by computing the entrance and exit points of each ray through the primitives.
-3D Convex Splatting (3DCS) [5] was inspired by the same shortcoming of GS: very large numbers of primitives are needed to model hard edges and flat surfaces due to the diffuseness of the Gaussians.
-As the name suggests, the scene is represented by a set of polyhedral convexes which undergo volumetric rendering, pruning and splitting operations.
-The concepts of smoothness and sharpness introduced by 3DCS have inspired our diffuseness (see Section 4).
-Radiant Foam [26] is both efficient, like GS, but also enables modeling light transport phenomena like reflection and refraction.
-The method takes its name from the representation which is a dense Voronoi tessellation of 3D space.
-The positions of Voronoi sites are differentiable after rendering, allowing iterative optimization.
-Since this is the only representation with connected elements, non-differentiable operations, e.g.
-pruning, have to be implemented carefully.
-All methods in this paragraph rely on different forms of piecewise constant volumetric elements with planar faces.
-Computing the intersections of each primitive with a ray requires multiple ray-triangle intersections, and is thus more expensive than our method.
-A few common themes emerge by analyzing the above methods.
-Most advocate the use of bounded primitives to enhance their ability to represent sharp edges and flat surfaces with small numbers of primitives.
-Our primitives are bounded but have diffuse boundaries to facilitate optimization.
-Like SolidGS [20], the opacity of our triangles remains approximately constant inside the triangle leading to sharper synthesized views.
-Several of the non-Gaussian primitives [5, 23, 25] leverage the existing software and hardware infrastructure by proposing representations with triangular faces.
-3 Background
-Before discussing our optimization framework, we first introduce the necessary background geometry used to parameterize a ray-triangle intersection in barycentric coordinates.
-The barycentric coordinates of a point on a triangle $\lambda=[\lambda^{0},\lambda^{1},\lambda^{2}]^{T}$ are one form of expressing the coordinates of the point as a linear combination of the coordinates of the three vertices, $[V^{0},V^{1},V^{2}]$.
-This parameterization is important for graphics applications to be able to efficiently rasterize triangles onto the screen space and interpolate the color from each vertex for a pixel.
-Throughout this paper, we follow the assumption that $\lambda^{j}\in\mathbb{R}^{+}$ corresponds to vertex $V^{j}\in\mathbb{R}^{3}$ such that $\sum_{j=1}^{3}\lambda^{j}=1$.
-Given a pixel, we compute the ray-triangle intersection,
-$P=C+(\widehat{r}d)$ (1)
-where $C\in\mathbb{R}^{3}$ is the camera center, $\widehat{r}\in\mathbb{R}^{3}$ is the unit vector for the ray through pixel $i$, and $d\in\mathbb{R}$ is the depth along the ray from the camera center to the intersection point, computed as follows,
-$d=\frac{\widehat{n}\cdot\overrightarrow{CB}}{\widehat{n}\cdot\widehat{r}}$ (2)
-where $B\in\mathbb{R}^{3}$ is the barycenter of the triangle and $\widehat{n}\in\mathbb{R}^{3}$ is its normal.
-We compute the barycentric coordinates for point $P$ using the triple products between the normal, a triangle edge, and the vector from each vertex to the point,
-$\lambda=\frac{1}{\widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}V_{2}})}\begin{bmatrix}\widehat{n}\cdot(\overrightarrow{V_{1}V_{2}}\times\overrightarrow{V_{1}P})\\ \widehat{n}\cdot(\overrightarrow{V_{2}V_{0}}\times\overrightarrow{V_{2}P})\\ \widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}P})\end{bmatrix}$ (3)
-Intuitively, the contribution of each vertex is proportional to the area of the sub-triangle formed by the intersection point $P$ and the other two vertices of the triangle.
-This weight becomes larger as $P$ approaches the vertex.
-4 Method
+The authors of SuGaR (Guédon & Lepetit, 2024) and PGSR (Chen et al., 2024a) devised similar techniques for aligning 2D Gaussian with the surfaces via regularization and a multi-view loss, respectively.
+Gaussian Surfels (Dai et al., 2024) rely on monocular surface normal estimates and a normal-depth consistency loss.
+Gaussian Opacity Fields (GOF) (Yu et al., 2024) extract surfaces as the zero-level set of 3D Gaussians.
+3D-Half-Gaussian Splatting (3D-HGS) (Li et al., 2024) enables the representation of perfectly planar surfaces and hard edges by attaching a splitting plane to each Gaussian, aligning it to the local surface, and setting the density of one of the half-spaces to 0.
 Figure 2: We begin with a set of sparse points.
@@
 Triangle edge neighbors are recomputed following adaptive densification.
-Given a set of images together with camera poses and a set of sparse points $\mathcal{S}$ computed via Structure-From-Motion [27], we construct an explicit scene representation using triangle primitives, the parameterization of which we discuss in Section 4.1.
+Representations based on NeRF and 3DGS have found great success, but have also inspired researchers to seek alternatives.
+NeuRBF (Chen et al., 2023) utilizes Radial Basis Functions (RBFs) to overcome limitations of NeRF due to the global nature of its MLP and features.
+GES (Hamdi et al., 2024) is based on an explicit representation which replaces the Gaussian kernel with a Generalized Exponential Function, overcoming the low-pass effect of the Gaussian and thus requiring fewer primitives to represent the scene.
+Similar approaches based on smooth, non-Gaussian kernels include DARB-Splatting (Arunan et al., 2025), SolidGS (Shen et al., 2024), Beyond Gaussians (Chen et al., 2024b) and Deformable Beta Splatting (Liu et al., 2025).
+Besides the methods that force their Gaussians to be planar (Huang et al., 2024; Guédon & Lepetit, 2024; Chen et al., 2024a), but still diffuse, there are others that represent the scene with collections of planar primitives.
+Zanjani et al.
+(2025) initially use Gaussian splats to model the scene and then merge them into 3D planes, which are abundant in indoor scenes.
+PlanarSplatting (Tan et al., 2025) is also designed for indoor scenes using planes, initialized via monocular depth estimation, as the only primitives.
+TRIPS (Franke et al., 2024) is based on the principle that point primitives can be rasterized into an image pyramid from which the appropriate layer can be selected according to the size of the projected point.
+Holes can be filled by a small network yielding accurate, crisp renderings.
+Triangle Splatting (Held et al., 2025a) uses triangles with diffuse edges as primitives, but does not support any mechanism for them to interact directly with each other.
+Non-planar primitives were introduced by BG-Triangle (Wu et al., 2025) which uses Bézier Gaussian triangles that are effective near boundaries, but comes at the cost of operating on complex, non-planar primitives that are hard to render.
+Quadratic Gaussian Splatting (Zhang et al., 2025) uses deformable quadratic surfaces as primitives and geodesic, instead of Euclidean, distance-based density distributions that adapt to the curvature of the primitives.
+Another class of methods rely on volumetric primitives.
+LinPrim (von Lützow & Nießner, 2025) is based on linear solid primitives with triangular faces and performs volumetric rendering by computing the entrance and exit points of each ray through the primitives.
+3D Convex Splatting (3DCS) (Held et al., 2025b) was inspired by the limitation of GS that requires very large numbers of primitives to model hard edges and flat surfaces due to the diffuseness of the Gaussians.
+As the name suggests, the scene is represented by a set of polyhedral convexes which undergo volumetric rendering, pruning and splitting operations.
+The concepts of smoothness and sharpness introduced by 3DCS have inspired our diffuseness (see Section 3).
+Radiant Foam (Govindarajan et al., 2025) enables modeling light transport phenomena, like reflection and refraction, by tessellating the space into Voronoi cells and iteratively optimizing the positions of the Voronoi vertices.
+All methods in this paragraph rely on different forms of volumetric elements with planar faces.
+Computing the intersections of each primitive with a ray requires multiple ray-triangle intersections, and is thus more expensive than RTS.
+A recent trend in the literature has been joint optimization of two representations: one for synthesizing novel views and one that is more faithful to the surfaces (Choi et al., 2024; Jiang et al., 2025).
+MILo (Guédon et al., 2025) maintains Gaussians that are alpha-blended for view synthesis and a watertight mesh without texture that captures the geometry of the scene.
+A few common themes emerge by analyzing the above methods.
+Most advocate the use of bounded primitives to enhance their ability to represent sharp edges and flat surfaces with small numbers of primitives.
+Our primitives are bounded but have diffuse boundaries to facilitate optimization.
+No other method, however, has a mechanism for direct inter-primitive communication like RTS.
+3 Method
+Figure 3: Triangle parameterization.
+Each triangle is parameterized by the incenter $\mu$, and three scale values $\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$.
+These parameters, along with the rotation matrix $R$, define the coordinates of each vertex $V^{j}$.
+Given a set of images together with camera poses and a set of sparse points $\mathcal{S}$ computed via Structure-From-Motion (Schönberger & Frahm, 2016), we construct an explicit scene representation using triangle primitives, the parameterization and initiliazation of which we discuss in Sections 3.1 and 3.2, respectively.
+The triangles are endowed with diffuse boundaries, similar to 3DCS (Held et al., 2025b) (Section 3.3).
 To model the surfaces in the scene, we render into each camera the image, depth, and normal map of the triangles through alpha-blending (see Section 3.4).
@@
 In our representation, primitives maintain soft connectivity with their neighbors, discussed in Section 3.5, for which we compute additional loss directly over the connection orientations (see Section 4).
-Similar to previous works [2, 3, 4, 1, 13], we develop an adaptive densification strategy in order to facilitate the addition and removal of primitives in the scene (see Section 4.6).
+Similar to previous works (Kerbl et al., 2023; Huang et al., 2024; Guédon & Lepetit, 2024; Chen et al., 2024a; Yu et al., 2024), we develop a strategy for adaptive density control in order to facilitate the addition and removal of primitives in the scene (see Section 3.6).
 We show an overview of our optimization algorithm in Fig.
@@
 $t_{n}=\{\mu,\Delta,s,R,\alpha,\sigma\}$ (1)
-where $\mu$ is the triangle reference point, $\Delta=\begin{bmatrix}\delta^{0}&\delta^{1}&\delta^{2}\end{bmatrix}$ are the per-vertex Spherical Harmonics, $s=\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$ are the vertex scales, $R$ is the $3\times 3$ rotation matrix, $\alpha$ is the opacity, and $\sigma$ is the diffuse scalar, discussed in Section 4.3.
+where $\mu$ is the incenter of the triangle, $\Delta=\begin{bmatrix}\delta^{0}&\delta^{1}&\delta^{2}\end{bmatrix}$ are the per-vertex Spherical Harmonics, $s=\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$ are the vertex scales, $R$ is the $3\times 3$ rotation matrix, $\alpha$ is the opacity, and $\sigma$ is the diffuse scalar, discussed in Section 3.3.
 3.2 Initialization
-The coordinates of the sparse SfM points are used as the initialization for the triangle reference points, and the point colors as the initialization for the zero-component of the Spherical Harmonics for all three vertices.
+The coordinates of each sparse SfM point are used as the initialization for the incenter, and the point colors as the initialization for the zero-component of the Spherical Harmonics for all three vertices, which are then optimized separately.
 Similar to previous works, the scale for our primitives is computed based on the average distance to the three nearest neighboring points.
-Each triangle starts as an equilateral triangle, using the scale to parameterize the distance of each vertex from the triangle reference point $\mu$ (see Fig.
+Each triangle starts as equilateral, using the scale to parameterize the distance of each vertex from the incenter of the triangle $\mu$ along the bisectors of the angles (see Fig.
 3).
-Note that in this representation, the aforementioned triangle reference point $\mu$ is not the barycenter of the triangle; that is, $\mu$ is not guaranteed to be the mean of all three vertices, but instead represents the anchoring position of the triangle.
 The diffuse scalar $\sigma$ is also set as a function of the average distance to the three nearest neighboring points.
 The parameterization of the initial rotation matrix for each triangle is defined such that the $\overrightarrow{V^{2}V^{0}}$ edge and the triangle normal $\widehat{n}$ are used as the $x$- and $z$- axis, respectively.
-We then initialize each triangle by applying a random rotation.
-Figure 3: Triangle parameterization.
-Each triangle is parameterized by a reference point $\mu$, and three scale values $\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$.
-These parameters, along with the rotation matrix $R$, define the coordinates of each vertex $V^{j}$.
+We then initialize each triangle by applying a random rotation to this matrix.
 3.3 Diffuse Primitive Boundaries
@@
 It is challenging in practice to optimize a scene with fully opaque primitives that do not smoothly blend with one another.
-It is therefore ideal to make the triangle primitives diffuse near the edges.
-Taking inspiration from 3DCS [5], we parameterize the response from each triangle as a function of the signed distance from the ray-triangle intersection to the nearest edge,
+It is therefore desirable to make the triangle primitives diffuse near the edges.
+Taking inspiration from 3DCS (Held et al., 2025b), we parameterize the response from each triangle as a function of the signed distance from the ray-triangle intersection to the nearest edge,
 $w_{\sigma}=\frac{1}{1+e^{(\sigma l)}}$ (2)
@@
 For the signed distance, $l<0$ occurs when the intersection point lies outside of the triangle boundary.
+This formulation slightly deviates from that of 3DCS, as the signed distance is computed directly, compared to their smooth approximation.
 Refer to caption Refer to caption Refer to caption
@@
 3.4 Rasterization
-With the barycentric coordinates of the ray-triangles intersections and diffuse weight computed, we interpolate the color for the current pixel w.r.t.
-the colors of each vertex weighted by the barycentric coordinates,
-$c_{n}=\begin{bmatrix}c^{0}&c^{1}&c^{2}\end{bmatrix}\lambda$ (6)
-where $c^{0}$, $c^{1}$, and $c^{2}\in\mathbb{R}^{3}$ are computed from the SH components of each vertex, and $\lambda$ are the barycentric coordinates of the intersection point.
-We aggregate the contribution to the current output pixel, $i$, from each intersected primitive,
-$c_{i}=\sum w_{n}c_{n}$ (7)
-where $w_{n}=\alpha w_{\sigma}T$, and $T=\Pi_{j=1}^{i-1}(1-\alpha_{j})$ is the transmittance for the current primitive [1].
-Previous works [2, 4, 13] provide two methods for rendering per-pixel depth; (1) computing the average weighted intersection depth of all traversed primitives (mean depth), and (2) using only the depth of the primitive that causes the transmittance $T$ to exceed $0.5$ (median depth).
+Throughout this paper, we use barycentric coordinates, $\lambda=[\lambda^{0},\lambda^{1},\lambda^{2}]^{T}$, to rasterize triangles.
+Please see Section S.1 for more details.
+With the barycentric coordinates of the ray-triangle intersections and the diffuse weight computed, we interpolate the color for the current pixel w.r.t.
+the colors of each vertex weighted by the barycentric coordinates, $c_{n}=\begin{bmatrix}c^{0}&c^{1}&c^{2}\end{bmatrix}\lambda$, where $c^{0}$, $c^{1}$, and $c^{2}\in\mathbb{R}^{3}$ are computed from the SH components of each vertex, and $\lambda$ are the barycentric coordinates of the intersection point.
+We aggregate the contribution to the current output pixel, $i$, from each intersected primitive, $c_{i}=\sum w_{n}c_{n}$, where $w_{n}=\alpha w_{\sigma}T$, and $T=\Pi_{j=1}^{i-1}(1-\alpha_{j})$ is the transmittance for the current primitive (Kerbl et al., 2023).
+Previous works (Huang et al., 2024; Chen et al., 2024a; Yu et al., 2024) provide two methods for rendering per-pixel depth; (1) computing the average weighted intersection depth of all traversed primitives (mean depth), and (2) using only the depth of the primitive that causes the transmittance $T$ to exceed $0.5$ (median depth).
 In our work, we use median depth.
-Using the average weighted depth encourages the formation of many semi-transparent layers of primitives.
+Using the mean depth encourages the formation of many translucent layers of primitives.
 Rendering median depth removes this blending and helps guide the formation of surfaces.
 We directly use $d$ from Eq.
-2 for the depth of each intersection.
-For each triangle, we compute the surface normal as the cross-product between two edges,
-$\widehat{n}_{n}=\frac{\overrightarrow{V^{0}V^{2}}\times\overrightarrow{V^{0}V^{1}}}{||\overrightarrow{V^{0}V^{2}}\times\overrightarrow{V^{0}V^{1}}||_{2}}$ (8)
-Unlike the rendered depth, to render per-pixel surface normals, we follow previous work [2, 4] and alpha-blend all the surface normals of all intersected primitive.
+S.2 for the depth of each intersection.
+Additionally, we compute the surface normal, $\widehat{n}_{n}$, for each triangle as the cross-product between two edges.
+Unlike the rendered depth, to render per-pixel surface normals, we follow previous work (Huang et al., 2024; Chen et al., 2024a) and alpha-blend all the surface normals of all intersected primitives.
 The intuition is that through a normal consistency loss (see Section 4), the blended normals must align with the normals computed from the median depth map.
 This encourages all the intersected triangles to align (and eventually collapse onto) the median surface.
-Thus, we render per-pixel surface normals through alpha-blending,
-$\overrightarrow{n}_{i}=\sum w_{n}\widehat{n}_{n}$ (9)
+Thus, we render per-pixel surface normals through alpha-blending, $\overrightarrow{n_{i}}=\sum w_{n}\widehat{n}_{n}$.
+Figure 5: Examples of edge association for connecting nearby triangles. For simplicity, it is assumed the triangles are co-planar in this figure.
 3.5 Connectivity
-To enable soft connectivity among primitives and encourage flat, connected surfaces, we add connectivity and surface normal terms to the optimization objective.
+To enable soft connectivity among primitives and encourage flat, connected surfaces, we add a connectivity term to the optimization objective.
 For each triangle edge, we assign, and periodically update, a connection to the nearest neighboring triangle edge.
-Taking inspiration from the energy functions introduced in 3D scene flow estimation [28], the loss of the connectivity increases according to the distance between their vertices and the inner product between their normals, discussed in Section 5.
-In order to prevent primitives from undesired connections, edge connection assignment must pass direction and orientation criteria, shown in Fig.
-5.
-Simply, we must ensure that the neighboring triangle edge is in the correct direction and is oriented appropriately in order to connect with the current edge.
-We compute connectivity assignment and activate the connectivity losses after several epochs.
-Refer to caption Refer to caption
-(a) Direction Constraint (b) Orientation Constraint
-Refer to caption Refer to caption
-(c) $\tau$ (d) $\rho$
-Figure 5: Above we show connectivity constraints between two triangles. For simplicity, it is assumed the triangles are co-planar.
-Naively connecting with the nearest edge without any consideration of relative position may lead to connections that cross triangles and could cause them to flip.
-To prevent such connections, we compute three unit vectors: (1) the unit vector orthogonal to the current triangle edge and the current triangle normal (pointing outside the triangle), (2) the unit vector orthogonal to the neighboring triangle edge and the neighboring triangle normal (pointing outside the triangle), and (3) the connecting vector from the center of the current triangle edge to the center of the neighboring triangle edge (normalized to be unit length).
-To measure the relative position between the triangles, we take the inner product between these unit vectors.
-The direction criterion is measured as the inner product between the current triangle edge vector and the connecting vector, shown in Fig.
-5(c).
-The orientation criterion is measured as the inner product between the current triangle edge vector and the neighboring triangle edge vector, shown in Fig.
-5(d).
-We apply thresholds, $\rho$ and $\tau$ as shown in the figure, to these two values to prevent any unwanted connectivity from forming.
-Intuitively, the direction criterion ensures the neighboring edge is in the correct direction w.r.t.
-the current triangle.
-Not satisfying this criterion would cause the current triangle to flip in order to reduce the connectivity loss.
-The orientation criterion ensures the edges are facing each other and the triangles have relatively similar surface normals.
-Not upholding this criterion would cause the neighboring triangle to flip in order to reduce the connectivity loss.
-4.6 Adaptive Densification
-Figure 6: During densification, large triangles are split into four sub-triangles with interpolated colors per-vertex.
-We perform adaptive density control through the process of cloning and splitting triangles.
-Previous works [2, 1, 4, 3] perform the cloning and splitting procedures by duplicating primitives conditioned on scale and position gradients.
+Taking inspiration from the energy functions introduced in 3D scene flow estimation (Vogel et al., 2015), the connectivity term of the loss increases according to the distance between their vertices and the inner product between their normals, discussed in Section 4.
+Naively connecting with the nearest edge without considering the relative orientations may cause connections that would require large changes in rotation to either triangle, leading to undesirable behaviors during optimization.
+To prevent this behavior, connections are only established if the triangle edges are "facing" each other.
+In practice, we use the inner product between the unit vectors orthogonal to the triangle edges (in the plane of each triangle) as the criterion for establishing connections.
+We provide an example in Fig.
+5 where a connection to the highlighted edge of triangle D is valid, while the other two connections would cause large rotations.
+3.6 Adaptive Density Control
+We perform adaptive density control through the process of cloning, splitting and pruning triangles.
+Previous works (Kerbl et al., 2023; Huang et al., 2024; Guédon & Lepetit, 2024; Chen et al., 2024a) perform the cloning and splitting procedures by duplicating primitives conditioned on scale and position gradients.
 In order to properly split large triangles, we must split them into four sub-triangles.
@@
 6 for a visualization of the splitting procedure.
+Figure 6: Triangles are split with interpolated colors per-vertex.
 Similar to previous work, we directly clone small primitives selected for densification.
-Much like 2DGS [2], our method does not directly rely on the gradient of the projected 2D primitive center.
-Instead of computing an approximation via projecting 3D gradients into the camera plane [2], we directly condition densification on the magnitude of the reference point gradient $\nabla\mu$ in 3D.
+Much like 2DGS (Huang et al., 2024), our method does not directly rely on the gradient of the projected 2D primitive center.
+Instead of computing an approximation via projecting 3D gradients into the camera plane (Huang et al., 2024), we directly condition densification on the magnitude of the incenter gradient $\nabla\mu$ in 3D.
+To remove uninformative triangles, we prune primitives that meet the following criteria: (i) triangles that are transparent ($\alpha<0.05$), (ii) triangles with one or fewer edge connections, (iii) triangles that do not intersect the camera frustum of at least three views after each epoch.
 4 Optimization
-Our objective function comprises terms computed in 2D, on the images planes, and in 3D, on the triangles.
-5.1 Rendering Losses
-Following previous work [1, 2, 3, 4, 13], we compute the SSIM between the rendered and input images,
-$\mathcal{L}_{ssim}=\frac{1}{N}\sum_{i,j}(1-\gamma)L_{1}+\gamma L_{ssim}$ (10)
-To help locally align the triangles with the rendered surface, we apply a normal consistency loss [2].
-This encourages alignment between the rendered normal $\widehat{n}$ with the normal computed from the rendered depth map $\widehat{n}_{d}$,
-$\mathcal{L}_{norm}=\frac{1}{N}\sum_{i,j}(1-\widehat{n}^{T}\widehat{n}_{d})$ (11)
-From the unsupervised depth estimation literature [29, 30, 31, 32, 33], we adopt a smoothness term on the rendered depth map conditioned on the gradient of the input image.
-This penalizes large gradients in the rendered depth map where we have small gradients in the input image,
-$\mathcal{L}_{smooth}=\frac{1}{N}\sum_{i,j}||\partial_{x}D_{i,j}||e^{||\partial_{x}I_{i,j}||}+||\partial_{y}D_{i,j}||e^{||\partial_{y}I_{i,j}||}$ (12)
-5.2 Scene Loss
-To encourage connectivity between primitives, we penalize the mean of the $L2$ distance between the connected vertex pairs of neighboring triangle edges,
-$\mathcal{L}_{conn}=\sum_{a\in\Omega}\frac{1}{2}(||V_{a}^{1}-V_{b}^{1}||_{2}+||V_{a}^{2}-V_{b}^{2}||_{2})+(1-\widehat{n}_{a}^{T}\widehat{n}_{b})$ (13)
-where $V_{a}^{j}$ and $V_{b}^{j}$ are the $j^{th}$ vertex pair of connected edges $a$ and $b$, respectively.
-$\widehat{n}_{a}$ and $\widehat{n}_{b}$ are the normals of the connecting triangles and $\Omega$ is the set of all triangles that intersect the current camera frustum.
-Activating this loss for invisible triangles without rendering losses leads to oversmoothing.
-The final objective is a weighted summation of the above,
-$\mathcal{L}=\omega_{0}\mathcal{L}_{ssim}+\omega_{1}\mathcal{L}_{norm}+\omega_{2}\mathcal{L}_{smooth}+\omega_{3}\mathcal{L}_{conn}$ (14)
-6 Experiments
-Table 1: Chamfer distance evaluation on scenes from DTU [6].
+Our objective function comprises terms computed in 2D (on the images plane), and in 3D (on the triangles).
+Table 1: Chamfer distance evaluation on scenes from DTU (Aanæs et al., 2016).
 Following previous literature, we average the accuracy and completeness Chamfer distances on the widely used evaluation set.
 Chamfer distances are measured in millimeters.
+The best results are in boldface and the second best are underlined.
 Method DTU
-24 37 40 55 63 65 69 83 97 105 106 110 114 118 122 Mean (mm)$\downarrow$
-3DGS [1] 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96
-SuGaR [3] 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
-2DGS [2] 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
-GOF [13] 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
-PGSR [4] 0.36 0.57 0.38 0.33 0.78 0.58 0.50 1.08 0.63 0.59 0.46 0.54 0.30 0.38 0.34 0.52
-RTS 0.48 0.59 0.79 0.39 0.58 0.92 0.77 0.94 0.77 0.37 0.79 1.12 0.37 0.71 0.49 0.67
-6.1 Implementation Details
-We implement the majority of our RTS framework in Python using PyTorch [34].
-For rasterization, we develop custom CUDA kernels for both the forward and backward pass.
-We run all our experiments on a single NVIDIA RTX A6000.
-In all of our experiments, we choose $\tau=0.0$ and $\rho=0.0$ for the threshold values on the connectivity constraints.
-For our loss weights, we choose $\omega=\begin{bmatrix}1.0&0.05&0.8&10.0\end{bmatrix}$ empirically.
-Following state-of-the-art Multi-View Stereo methods [35, 36, 37, 38, 39, 40], we directly generate a final 3D point cloud from the rendered depth maps.
-To generate our final point cloud, we use heuristic filtering, similar to the post-processing presented in GBiNet [40].
-For each rendered depth map, we project each depth estimate into neighboring views, using the depth estimates in each view at the projected pixel location to reproject back to the reference view, measuring the pixel reprojection error and filtering out estimates whose error is above a threshold.
-6.2 Evaluation
-We test our framework on DTU [6], which is an indoor dataset that contains images of 124 scenes taken from a camera mounted on an industrial robot arm.
-All scenes share the same camera trajectories, with ground-truth point clouds captured via structured light.
-We evaluate our new approach on the DTU dataset and record the Chamfer distance in Table 1.
-We show competitive results alongside the leading state-of-the-art SGS methods.
-Across all scenes in the test set, we report highly competitive metrics against the current state of the art.
-Notably, our method seems particularly sensitive to specific scenes with texture-less and non-Lambertian surfaces.
-Qualitative results can be seen in Fig.
-7.
+[…]
+3DGS (Kerbl et al., 2023) 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50 1.96
+SuGaR (Guédon & Lepetit, 2024) 1.47 1.33 1.13 0.61 2.25 1.71 1.15 1.63 1.62 1.07 0.79 2.45 0.98 0.88 0.79 1.33
+2DGS (Huang et al., 2024) 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
+GOF (Yu et al., 2024) 0.50 0.82 0.37 0.37 1.12 0.74 0.73 1.18 1.29 0.68 0.77 0.90 0.42 0.66 0.49 0.74
+Gaussian Surfels (Dai et al., 2024) 0.66 0.93 0.54 0.41 1.06 1.14 0.85 1.29 1.53 0.79 0.82 1.58 0.45 0.66 0.53 0.88
+PGSR (Chen et al., 2024a) 0.36 0.57 0.38 0.33 0.78 0.58 0.50 1.08 0.63 0.59 0.46 0.54 0.30 0.38 0.34 0.52
+TriangleSplatting (Held et al., 2025a) 0.98 1.07 1.07 0.51 1.67 1.44 1.17 1.32 1.75 0.98 0.96 1.11 0.56 0.93 0.72 1.06
+RTS 0.42 0.61 0.74 0.39 0.53 0.86 0.70 0.84 0.72 0.37 0.68 0.87 0.34 0.58 0.44 0.61
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Ground Truth Rendered Image Rendered Depth Rendered Normal Point Cloud
-Figure 7: Qualitative results on several scenes from the DTU dataset [6].
-We measure the quality of novel view synthesis for our method on the DTU dataset using standard metrics of PSNR, SSIM, and LPIPS [41] in Table 2.
-Our method produces comparable results, with a slight reduction in PSNR.
-As von Lützow and Nießner [25] note, using primitives with explicit boundaries can begin to introduce hard edges in regions with poor visibility.
-While the diffuse boundaries of our triangles helps prevent much of this behavior, its effects are noticeable in some of the reconstructed images.
-Table 2: Photometric evaluation on several scenes from the DTU [6] dataset. We record the per-scene SSIM, PSNR, and LPIPS metrics on the widely used test set.
-Method DTU
-SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$
-SuGaR [3] - 34.57 -
-2DGS [2] 0.939 34.33 0.169
-PGSR [4] 0.917 33.18 0.251
-RTS 0.918 31.68 0.213
+Figure 7: Qualitative results on the DTU dataset (Aanæs et al., 2016).
+RTS estimates high-quality geometry, maintaining thin structures, such as the carrot noses and branches (middle row) and the book edges (bottom row), while effectively modeling smooth surfaces, such as the Buddha (top row).
+4.1 Rendering Losses
+Following previous work (Kerbl et al., 2023; Huang et al., 2024), we compute the SSIM between the rendered and input images, $\mathcal{L}_{ssim}$, and apply a normal consistency loss (Huang et al., 2024), $\mathcal{L}_{norm}$, to help locally align the triangles with the rendered surface.
+Please see Section S.3 the supplement.
+From the unsupervised depth estimation literature (Chang et al., 2022; Godard et al., 2017; Mahjourian et al., 2018), we adopt a smoothness term on the rendered depth map conditioned on the gradient of the input image.
+This penalizes large gradients in the rendered depth map where we have small gradients in the input image,
+[…]
+4.2 Scene Loss
+To encourage connectivity between primitives, we penalize the mean of the $L2$ distance between the connected vertex pairs of neighboring triangle edges,
+[…]
+where $V_{a}^{j}$ and $V_{b}^{j}$ are the $j^{th}$ vertex pair of connected edges $a$ and $b$, respectively.
+$\widehat{n}_{a}$ and $\widehat{n}_{b}$ are the normals of the connecting triangles and $\Omega$ is the set of all triangles that intersect the current camera frustum.
+Applying this loss to invisible triangles without rendering losses leads to over-smoothing.
+The final objective is a weighted summation of all terms:
+$\mathcal{L}=\omega_{0}\mathcal{L}_{ssim}+\omega_{1}\mathcal{L}_{norm}+\omega_{2}\mathcal{L}_{smooth}+\omega_{3}\mathcal{L}_{conn}$ (5)
+5 Experiments
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Ground Truth RTS 2DGS (Huang et al., 2024)
+Figure 8: Qualitative comparison between RTS and 2DGS (Huang et al., 2024) on; Top - scan024 from the DTU dataset (Aanæs et al., 2016), Bottom - kitchen from the mip-NeRF 360 dataset (Barron et al., 2022).
+RTS is substantially more precise at estimating the geometry at discontinuities and rendering fine details, shown in the areas marked by the blue rectangles.
+5.1 Implementation Details
+We implement the majority of our RTS framework in Python using PyTorch (Paszke et al., 2019).
+For rasterization, we develop custom CUDA kernels for both the forward and backward pass.
+We run all our experiments on a single NVIDIA RTX A6000.
+For our loss weights, we choose $\omega=\begin{bmatrix}1.0&0.05&50.0&1000.0\end{bmatrix}$ empirically.
+Following state-of-the-art Multi-View Stereo methods (Yao et al., 2018; Yang et al., 2022; Mi et al., 2022), we directly generate a 3D point cloud for geometric evaluation from the rendered depth maps without performing any TSDF fusion.
+To generate each point cloud, we use simple heuristic filtering on each depth map, similar to the post-processing presented in GBiNet (Mi et al., 2022).
+For each depth map, we measure the reprojection error of the depth values at every pixel using neighboring views and filter pixels based on this error.
+All depth estimates with a low reprojection error are back-projected to 3D points, forming the combined point cloud.
+Table 2: Novel View Synthesis on all scenes from the mip-NeRF 360 dataset (Barron et al., 2022).
+Method Outdoor Indoor
+[…]
+NeRF (Mildenhall et al., 2020) 21.46 0.458 0.515 26.84 0.790 0.370
+Deep Blending (Hedman et al., 2018) 21.54 0.524 0.364 26.40 0.844 0.261
+Instant NGP (Müller et al., 2022) 22.90 0.566 0.371 29.15 0.880 0.216
+MipNeRF360 (Barron et al., 2022) 24.47 0.691 0.283 31.72 0.917 0.180
+SuGaR (Guédon & Lepetit, 2024) 22.93 0.629 0.356 29.43 0.906 0.225
+3DGS (Kerbl et al., 2023) 24.64 0.731 0.234 30.41 0.920 0.189
+2DGS (Huang et al., 2024) 24.34 0.717 0.246 30.40 0.916 0.195
+GOF (Yu et al., 2024) 24.82 0.750 0.202 30.79 0.924 0.184
+PGSR (Chen et al., 2024a) 24.76 0.752 0.203 30.36 0.934 0.147
+3DCS (Held et al., 2025b) 24.07 0.700 0.238 31.33 0.927 0.166
+TriangleSplatting (Held et al., 2025a) 24.27 0.722 0.217 30.80 0.928 0.160
+RTS 21.41 0.657 0.349 30.28 0.921 0.130
+5.2 Evaluation
+We test our framework on the DTU dataset (Aanæs et al., 2016), an indoor dataset that contains images of 124 scenes taken from a camera mounted on an industrial robot arm.
+All scenes share the same camera trajectories, with ground-truth point clouds captured via structured light.
+We evaluate our new approach on the DTU dataset and record the Chamfer distance in Table 1.
+We show competitive results alongside the leading state-of-the-art planar GS methods.
+Across the test set, RTS is the most geometrically accurate in several scenes and second most overall.
+While PGSR demonstrates impressive reconstruction results, the algorithm utilizes a full suite of multi-view objective functions that significantly improve the geometric reconstruction quality.
+We provide qualitative results on the DTU dataset in Fig.
+7, showing visualizations of the rendered images, depth maps, normal maps, and final point clouds for three scenes.
+In Fig.
+8 (top), we provide a comparison of depth map renderings between RTS and 2DGS (Huang et al., 2024).
+RTS is able to reconstruct fine details on the surfaces of objects that are typically blurred with Gaussian representations.
+Table 3: Ablation study on the contribution of the 3D loss terms using the entire DTU evaluation set.
+Here, $\lambda_{c}$ is the loss weight for the connectivity term, $L_{conn}$, and $\lambda_{s}$ is the loss weight for the depth smoothness term $L_{smooth}$.
+The best results are in boldface and the worst are underlined.
+Method Acc.($m$) $\downarrow$ Comp.($m$) $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Primitives (K) $\downarrow$
+w/o $\mathcal{L}_{smooth}$ & $\mathcal{L}_{conn}$ 0.66 0.68 0.912 30.58 0.227 218
+w/o $\mathcal{L}_{smooth}$ 0.65 0.71 0.921 32.13 0.212 244
+w/o $\mathcal{L}_{conn}$ 0.67 0.69 0.908 30.15 0.231 224
+full ($\lambda_{c}=10.0$, $\lambda_{s}=0.8$) 0.64 0.70 0.918 31.68 0.213 249
+full ($\lambda_{c}=300.0$, $\lambda_{s}=20.0$) 0.61 0.63 0.910 30.87 0.232 244
+full ($\lambda_{c}=1000.0$, $\lambda_{s}=50.0$) 0.59 0.62 0.909 30.57 0.232 297
+We show additional results on the mip-NeRF 360 dataset (Barron et al., 2022).
+Following the protocol specified by Barron et al.
+(2022), we separate the images in each scene, taking every eighth image as a test image and training on the remaining.
+As standard evaluation, we report PSNR, SSIM, and LPIPS (Zhang et al., 2018) metrics.
+We show results on all indoor and outdoor scenes in Table 2, as well as per-scene results in Table S.1.
+For indoor scenes, RTS shows results on par with the state-of-the-art methods, having the leading LPIPS score, the fourth overall SSIM score, and a highly competitive PSNR score among all listed methods.
+On two outdoor scenes (treehill and flowers), RTS is limited in reconstructing extremely distant background foliage, impacting the overall metrics (please see Fig.
+S.3 for qualitative results on these scenes).
+In Fig.
+8 (bottom), we provide a comparison of novel view synthesis between RTS and 2DGS.
+RTS is able to render fine textures in low visibility regions in scenes as opposed to blurring with Gaussian representations.
+5.3 Ablations
+We show an ablation evaluating the contributions of proposed loss terms in Table 3.
+Removing the soft connectivity leads to a decrease in overall accuracy of the output models, while removing the depth supervision negatively affects the completeness.
+The two supervision signals complement each other, and we show that increasing the loss weights allows for tuning the framework for either better novel view synthesis or geometry.
+Additionally, the results corresponding to the changes in magnitude of the loss weights demonstrates the stability of RTS to changes in hyper-parameters.
+To evaluate geometry, all previous planar GS algorithms utilize a GT foreground-background segmentation mask when generating the final models.
+Using this mask to generate the final models for evaluation removes the effects of floaters and inaccurate estimation near the surfaces being evaluated.
+To portray a more grounded evaluation of geometry, we provide an ablation study in Table S.2 in which we compute Chamfer distance of models without the use of any GT masks.
 6 Limitations & Conclusions
-Our geometric loss terms, much like 2DGS, involves a trade-off between model reconstruction quality and novel-view synthesis.
-Adhering more strictly to the geometric losses in our framework can lead to over-smoothing in some areas, and at times can prohibit the movement of primitives, as triangles are more strongly attracted to each other, leaving parts of the scene under-reconstructed.
-Furthermore, due to the periodic nearest-neighbors search, the run-time for our method increases with the number of primitives.
-Because of this, it can become difficult to reconstruct very large scenes.
 In this work, we introduce a new scene representation, namely Radiant Triangle Soup (RTS).
 To the best of our knowledge, we are the first to introduce explicit 3D forces between primitives in a splatting framework, helping to coordinate the positioning of primitives to directly form surfaces.
-The introduction of Triangle Soup as the scene representation for radiance fields is amenable to future work in surface optimization and can be augmented to include more extensive forms of triangle connectivity and mesh evolution.
-After some modifications, our framework can also be used for optimization over watertight meshes.
-In future work, we plan to extend RTS with modified primitive connectivity strategies and mesh post-processing refinement.
-Supplementary Material
-S.1 Implementation Details
+Modifying the weights of these forces allows for tuning between 3D reconstruction quality and novel-view synthesis quality.
+The main limitation of our current algorithm is its inability to extract watertight meshes.
+Furthermore, due to the periodic nearest-neighbors search, there is a minor increase in run-time proportional to the number of primitives.
+Please see Section S.2 for more details.
+The introduction of Triangle Soup as the underlying representation for Radiance Fields is amenable to future work in surface optimization.
+We plan to extend RTS with modified primitive connectivity strategies and perform optimization over watertight meshes.
+Supplemental Material
+Here we include additional material on background geometry used for computing barycentric coordinates, further implementation details, and additional results.
+S.1 Background
+In this section, we introduce the necessary background geometry used to parameterize a ray-triangle intersection in barycentric coordinates.
+The barycentric coordinates of a point on a triangle $\lambda=[\lambda^{0},\lambda^{1},\lambda^{2}]^{T}$, s.t.
+$\sum_{j=1}^{3}\lambda^{j}=1$, provide a means for expressing the coordinates of the point as a linear combination of the coordinates of the three vertices, $V^{0},V^{1},V^{2}]$.
+This parameterization is important for graphics applications to be able to efficiently rasterize triangles onto screen space and interpolate the color of a pixel from each vertex.
+In rendering, we compute the ray-triangle intersections using,
+$P=C+(\widehat{r}d)$ (S.1)
+where $C\in\mathbb{R}^{3}$ is the camera center, $\widehat{r}\in\mathbb{R}^{3}$ is the unit vector for the ray through pixel $i$, and $d\in\mathbb{R}$ is the depth along the ray from the camera center to the intersection point, computed as follows,
+$d=\frac{\widehat{n}\cdot\overrightarrow{CB}}{\widehat{n}\cdot\widehat{r}}$ (S.2)
+where $B\in\mathbb{R}^{3}$ is the barycenter of the triangle and $\widehat{n}\in\mathbb{R}^{3}$ is its normal.
+We compute the barycentric coordinates for point $P$ using the triple products between the normal, a triangle edge, and the vector from each vertex to the point,
+$\lambda=\frac{1}{\widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}V_{2}})}\begin{bmatrix}\widehat{n}\cdot(\overrightarrow{V_{1}V_{2}}\times\overrightarrow{V_{1}P})\\ \widehat{n}\cdot(\overrightarrow{V_{2}V_{0}}\times\overrightarrow{V_{2}P})\\ \widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}P})\end{bmatrix}$ (S.3)
+Intuitively, the contribution of each vertex is proportional to the area of the sub-triangle formed by the intersection point $P$ and the other two vertices of the triangle.
+This weight becomes larger as $P$ approaches the vertex.
+S.2 Implementation Details
 In this section, we describe our experimental setting and optimization parameters in detail.
-To begin optimization, similar to previous work [4, 2], all geometric supervision is disabled, with optimization only being guided initially by the SSIM loss $\mathcal{L}_{ssim}$.
+To begin optimization, similar to previous work (Huang et al., 2024; Chen et al., 2024a), all geometric supervision is disabled, with optimization only being guided initially by the SSIM loss $\mathcal{L}_{ssim}$.
 We enable the normal consistency loss $\mathcal{L}_{norm}$ at iteration $7,000$ and enable both the smoothness loss $\mathcal{L}_{smooth}$ and connectivity loss $\mathcal{L}_{conn}$ at iteration $10,000$, both chosen empirically.
@@
 Densification and pruning is run every $250$ iterations starting after iteration $2,000$.
-The maximum reference point gradient threshold for densification is set to $7.5e^{-5}$ in all experiments.
+The maximum incenter gradient threshold for densification is set to $7.5e^{-5}$ in all experiments.
 The learning rates for each respective parameter are set as follows:
@@
   • Opacity ($\alpha$): $5e^{-2}$
-  • Reference Point ($\mu$): $\begin{bmatrix}1.5e^{-4},&2e^{-6}\end{bmatrix}$
+  • Incenter ($\mu$): $\begin{bmatrix}1.5e^{-4},&2e^{-6}\end{bmatrix}$
   • Rotation ($R$): $1e^{-3}$
@@
   • Diffuse Scalar ($\sigma$): $1e^{-3}$
-where the Reference Point learning rate follows an exponential decay scheduler starting with $1.5e^{-4}$ and ending with $2e^{-6}$ conditioned on the total number of iterations.
+where the incenter learning rate follows an exponential decay scheduler starting with $1.5e^{-4}$ and ending with $2e^{-6}$ conditioned on the total number of iterations.
 We also tested a linear decay scheduler leading to similar results.
+Edge Connectivity Overhead
+To compute neighboring edge connections, we construct a single KD-Tree containing the midpoints of all triangle edges which is then queried for each edge once.
+The connected edge indices are the only structure that is stored during optimization.
+Since the parameters of the primitives are modified during optimization, we need to recompute the KD-Tree and connected edge indices every 250 iterations (aligned with Adaptive Density Control), which ultimately leads to a minimal overhead for moderate size scenes.
+To be concrete, building the KD-Tree and computing the neighboring indices takes on average 4 seconds for around 300,000 primitives, which is more than the average number of primitives needed to reconstruct the scenes for the DTU dataset.
+Since this operation only happens every 250 iterations starting after iteration 10,000 (when the connectivity loss is activated), the overhead of this operation only adds roughly 5 minutes to the optimization process, which is about a $6\%$ increase in runtime.
+Reconstructing larger scenes, or more precisely, scenes that require more primitives, will naturally demand a larger overhead.
+Run-time
 The run-times for our approach are roughly $1.5$ hours on DTU and $4.5$ hours on Mip-NeRF 360, with scenes from DTU and scenes from Mip-NeRF 360 having on average $249,941$ and $1,275,985$ primitives, respectively, using $0.5$ resolution for the DTU dataset and $0.25$ resolution for the Mip-NeRF 360 dataset.
-S.2 Ablation on Depth Rendering
-We show an ablation on the effects of using mean and median depth, in Table S.1.
-While using median depth has virtually no effect on the quality of the reconstructed images, there is a substantial effect on the accuracy and consistency of the rendered depth maps, and thus, the final geometry of the 3D model.
-The normal consistency loss $\mathcal{L}_{norm}$ is indifferent to surface layering when computed w.r.t.
-the blended mean depth and the blended mean normal.
-Using the median depth guides the median triangle to better align with the blended normals and vice versa.
-Table S.1: Comparison of the effects of mean vs median depth using the entire DTU evaluation set.
-Method Acc.(mm) $\downarrow$ Comp.(mm) $\downarrow$ PSNR $\uparrow$ Primitives $\downarrow$
-Mean Depth 1.23 1.36 31.69 228,000
-Median Depth 0.64 0.70 31.68 249,000
-Table S.2: Novel View Synthesis on all scenes from the Mip-NeRF 360 [42] dataset.
-Method Outdoor Indoor
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure S.1: Point cloud visualization for all scenes from the evaluation set from the DTU dataset(Aanæs et al., 2016).
+S.3 Loss Functions
+For completeness, we define the loss terms used in this paper that were introduced in previous work (Kerbl et al., 2023; Huang et al., 2024). The SSIM loss is computed as follows:
+$\mathcal{L}_{ssim}=\frac{1}{N}\sum_{i,j}(1-\gamma)L_{1}+\gamma L_{D\text{-}SSIM}$ (S.4)
+The normal consistency loss is computed as follows:
+$\mathcal{L}_{norm}=\frac{1}{N}\sum_{i,j}(1-\widehat{n}^{T}\widehat{n}_{d})$ (S.5)
+This encourages alignment between the rendered normal $\widehat{n}$ with the normal computed from the rendered depth map $\widehat{n}_{d}$,
+S.4 Additional Evaluations
+Table S.1: Novel View Synthesis per-scene on all scenes from the mip-NeRF 360 dataset (Barron et al., 2022).
+Scene SSIM PSNR LPIPS
+room 0.919 31.25 0.156
+counter 0.900 28.27 0.139
+kitchen 0.936 30.69 0.083
+bonsai 0.930 30.93 0.142
+bicycle 0.637 20.31 0.365
+flowers 0.617 19.99 0.351
+garden 0.742 24.89 0.250
+stump 0.658 23.25 0.410
+treehill 0.631 18.64 0.371
+In Table S.1, we show the per-scene novel view synthesis evaluation for RTS on the mip-NeRF 360 dataset (Barron et al., 2022) using the standard metrics of PSNR, SSIM, and LPIPS.
+As von Lützow & Nießner (2025) note, using primitives with explicit boundaries can begin to introduce hard edges in regions with poor visibility while smoother primitives degrade more gracefully.
+While the diffuse boundaries of our triangles help prevent much of this behavior, its effects are noticeable in some of the reconstructed images, especially in outdoor scenes with distant background foliage.
+See Fig.
+S.3
+As mentioned in Section 5.3, we compare the Chamfer distances of RTS and two competitive GS algorithms in Table S.2 with and without the use of the GT segmentation masks.
+While the distances increase for all methods, this experiment demonstrates how RTS is more effective at floater removal and background modeling.
+We provide additional renderings of novel views on challenging indoor and outdoor scenes with fine details and non-Lambertian surfaces.
+We show qualitative geometric reconstruction results on all scenes from the DTU dataset in Fig.
+S.1 and novel view synthesis results on all indoor and outdoor scenes from the Mip-NeRF 360 datasets in Fig.
+S.2 and Fig.
+S.3, respectively.
+Table S.2: Chamfer distance evaluation on scenes from DTU (Aanæs et al., 2016).
+Following previous literature, we average the accuracy and completeness Chamfer distances on the widely used evaluation set.
+Chamfer distances are measured in millimeters.
+Top: Chamfer distance evaluation using the GT mask.
+Bottom: Chamfer distance evaluation without using the GT mask.
+Method DTU
 […]
-NeRF [11] 21.46 0.458 0.515 26.84 0.790 0.370
-Deep Blending [43] 21.54 0.524 0.364 26.40 0.844 0.261
-Instant NGP [16] 22.90 0.566 0.371 29.15 0.880 0.216
-MipNeRF360 [42] 24.47 0.691 0.283 31.72 0.917 0.180
-SuGaR [3] 22.93 0.629 0.356 29.43 0.906 0.225
-3DGS [1] 24.64 0.731 0.234 30.41 0.920 0.189
-2DGS [2] 24.34 0.717 0.246 30.40 0.916 0.195
-GOF [13] 24.82 0.750 0.202 30.79 0.924 0.184
-PGSR [4] 24.76 0.752 0.203 30.36 0.934 0.147
-3DCS [5] 24.07 0.700 0.238 31.33 0.927 0.166
-RTS 21.41 0.657 0.349 30.28 0.921 0.130
-S.3 Additional Evaluations
-In this subsection, we show additional results on the Mip-NeRF 360 dataset [42].
-Following the protocol specified by the authors of Mip-NeRF 360, we separate the images in each scene, taking every eighth image as a test image and training on the remaining.
-As standard evaluation, we report PSNR, SSIM, and LPIPS [41] metrics.
-We show results on all indoor and outdoor scenes in Table S.2.
-For indoor scenes, RTS shows results on par with the state-of-the-art methods, having the leading LPIPS score, the fourth overall SSIM score, and a highly competitive PSNR score among all listed methods.
-On two outdoor scenes (treehill and flowers), RTS shows difficulty reconstructing extremely distant background foliage, impacting the overall metrics.
-We provide supplemental renderings of novel views on challenging indoor and outdoor scenes with fine details and non-Lambertian surfaces.
-Output renderings and 3D models for all scenes will be made publicly available to the research community if the paper is accepted.
-We show qualitative geometric reconstruction results on all scenes from the DTU dataset in Fig.
-S.2 and novel view synthesis results on all indoor and outdoor scenes from the Mip-NeRF 360 datasets in Fig.
-S.1 and Fig.
-S.3, respectively.
+2DGS (Huang et al., 2024) 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52 0.80
+PGSR (Chen et al., 2024a) 0.36 0.57 0.38 0.33 0.78 0.58 0.50 1.08 0.63 0.59 0.46 0.54 0.30 0.38 0.34 0.52
+RTS 0.42 0.61 0.74 0.39 0.53 0.86 0.70 0.84 0.72 0.37 0.68 0.87 0.34 0.58 0.44 0.61
+2DGS (Huang et al., 2024) [no mask] 1.22 1.69 0.88 0.43 0.96 0.77 0.85 1.23 1.87 1.69 0.91 2.01 0.82 0.82 1.07 1.15
+PGSR (Chen et al., 2024a) [no mask] 1.10 1.59 0.97 0.45 1.98 0.64 0.59 1.70 1.69 1.57 0.74 0.63 0.44 0.71 0.74 1.04
+RTS [no mask] 0.52 0.92 0.94 0.55 0.91 1.01 0.74 1.09 1.05 0.67 0.88 1.07 0.42 0.79 0.80 0.82
 Refer to caption Refer to caption Refer to caption
@@
 (a) GT Image (b) Rendered Image (c) Rendered Depth
-Figure S.1: Visualizations of all indoor scenes, (top-down) [bonsai, counter, kitchen, room], from the Mip-NeRF 360 dataset [42].
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure S.2: Point cloud visualization for all scenes from the evaluation set from the DTU dataset [6].
+Figure S.2: Visualizations of all indoor scenes, (top-down) [bonsai, counter, kitchen, room], from the mip-NeRF 360 dataset (Barron et al., 2022).
 Refer to caption Refer to caption Refer to caption
@@
 (a) GT Image (b) Rendered Image (c) Rendered Depth
-Figure S.3: Visualizations of all outdoor scenes, (top-down) [bicycle, flowers, garden, stump, treehill], from the Mip-NeRF 360 dataset [42].
+Figure S.3: Visualizations of all outdoor scenes, (top-down) [bicycle, flowers, garden, stump, treehill], from the Mip-NeRF 360 dataset (Barron et al., 2022).
+In the top-right section of the rendered image of treehill, RTS imprecisely approximates the appearance of the background foliage, while GS-based algorithms typically blur this region.